The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 13,750 $ 17,250 $ 31,000
Estimated variable manufacturing overhead per machine-hour $ 2.90 $ 3.70
Job P Job Q
Direct materials $ 28,000 $ 15,500
Direct labor cost $ 33,000 $ 13,500
Actual machine-hours used:
Molding 3,200 2,300
Fabrication 2,100 2,400
Total 5,300 4,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-9, assume that Sweeten Company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments and Job P included 20 units and Job Q included 30 units. For questions 10-15, assume that the company uses a plantwide predetermined overhead rate with machine-hours as the allocation base.

Foundational 2-11 (Algo)

11. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)
Answer #1

Predetermine overhead rate = (2500*2.9+1500*3.7+31000)/4000 = 10.95 per MH

11) Overhead applied

Job P Job Q
Overhead applied 5300*10.95 = 58035 4700*10.95 = 51465